NASA GOES SATOPS Morning Report: Friday, March 04, 2005

Status Report From: National Oceanic and Atmospheric Administration
Posted: Friday, March 4, 2005

DAY 63


OPS 54 Rev 07099 R+02 at HULA-B at 0441z on March 4: Received telemetry hits throughout entire support from rise to fade. HULA attempted to reissue a track order and this did not solve the problem. Never gained command access to the vehicle. Added up FBKS-C to recover data. USER declared a CAT 1 data loss on the smooth data due to timeline constraints (6 minutes late).

All DMSP operations were nominal over the past 24 hours.


1908z on March 3 Lost communication with Fairbanks comm. controllers FCOMM1 and FCOMM2. Had to reschedule a NOAA 17 Rev 13985 pass to the Rev 13986 but all data was recovered. Frame sync problems occurred throughout shift but the crews did a magnificent job recovering all recorded data. Resetting routeabout at SOCC appeared to fix the problem. Ground engineers are currently investigating.

All POES operations were nominal over the past 24 hours.


All other GOES operations were nominal over the past 24 hours.

Questions or comments to

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter